# ShapeLLM: Universal 3D Object Understanding for Embodied Interaction

### v2

@@ Abstract
This paper presents ShapeLLM, the first 3D Multimodal Large Language Model (LLM) designed for embodied interaction, exploring a universal 3D object understanding with 3D point clouds and languages. ShapeLLM is built upon an improved 3D encoder by extending ReCon [129] to ReCon++ that benefits from multi-view image distillation for enhanced geometry understanding. By utilizing ReCon++ as the 3D point cloud input encoder for LLMs, ShapeLLM is trained on constructed instruction-following data and tested on our newly human-curated evaluation benchmark, 3D MM-Vet. ReCon++ and ShapeLLM achieve state-of-the-art performance in 3D geometry understanding and language–unified 3D interaction tasks, such as embodied visual grounding.
Figure 1: Demonstrations of ShapeLLM. We present ShapeLLM, a multi-modal large language model designed for embodied scenes. ShapeLLM can generate accurate and prompt responses to user queries by leveraging comprehensive prior knowledge of the language. This figure demonstrates the powerful capabilities of ShapeLLM in planning & task decomposition, 3D vision question answering and embodied vision grounding, etc. The predicted bounding box is shown in the red box from three distinct perspectives.
@@ 1 Introduction
    Selective Multi-View Distillation  Interacting with objects typically necessitates an intricate 3D understanding that involves knowledge at various levels and granularities. For instance, a whole-part high-level semantic understanding is needed for interactions like opening a large cabinet, while detailed, high-resolution (i.e., low-level) semantics are crucial for smaller objects like manipulating a drawer handle [177]. However, existing works mainly distill single-view high-resolution object features from 2D foundation models [134], providing a complementary understanding [36, 129, 169]. The potential of multi-view images, which offer abundant multi-level features due to view variation and geometry consistency [9, 143, 79, 58, 101, 63], is often neglected. ShapeLLM extends ReCon [129] to ReCon++  as the 3D encoder by integrating multi-view distillation. To enable the model to selectively distill views that enhance optimization and generalization, inspired by Carion et al., ReCon++  is optimized through adaptive selective matching using the Hungarian algorithm [82].
@@
    3D Visual Instruction Tuning  Instruction tuning has been proven effective in improving LLMs’ alignment capability [117, 121]. To realize various 3D understanding tasks with a universal language interface, ShapeLLM is trained through instruction-following tuning on constructed language-output data. However, similar to 2D visual instruction tuning [93, 4], the data-dessert issue [36] is even worse since no object-level VQA data is available, unlike 2D [92]. To validate the efficacy of ShapeLLM, we first construct $\sim$45K instruction-following data using the advanced GPT-4V(ision) [115] on the processed Objaverse dataset [29] and 30K embodied part understanding data from GAPartNet [47] for supervised fine-tuning. Following MM-Vet [179], we further develop a novel evaluation benchmark named 3D MM-Vet. This benchmark is designed to assess the core vision-language capabilities, including embodied interaction in a 3D context, thereby stimulating future research. The 3D MM-Vet benchmark comprises 59 diverse Internet [^1] URL & License. 3D objects and 232 human-written question-answer pairs.
@@ 2 ShapeLLM
In this section, we first introduce the overall architecture of ShapeLLM. Then, we delve into two critical challenges faced in interactive 3D understanding: data dessert [36] and representation of 3D point clouds. We present the detailed design of our method to tackle these challenges, respectively.
@@ 2.1 Overall Architecture
Figure 3: Construction illustration of instruct-following data using GPT-4V [115]. Four perspective views are input into GPT-4V. In-context prompts focusing on different topics are explicitly incorporated to ensure data diversity.
2.2 How to alleviate interactive 3D understanding Data Dessert?
Most published 3D data is typically presented as 3D object-caption pairs, lacking an interactive style. Although a few concurrent works [166, 62] have attempted to construct interactive 3D understanding datasets, the questions-and-answers (Q&As) are primarily based on annotated captions, often providing a limited perspective without sufficient details. Additionally, those works have generally been limited to semantic understanding without considering embodied interaction. To address these limitations, our work constructs question-and-answer pairs based on multi-view images of a 3D object using GPT-4V(ision) [115]. For data diversity, we explicitly introduce six aspects as prompts, as illustrated Fig. 3. In the following, we provide the details about data collection and construction regarding general semantic understanding and embodied object understanding, respectively.
@@ 2.2 How to alleviate interactive 3D understanding Data Dessert?
General Semantic Understanding  This aims to enhance the model’s generalization abilities in visual recognition, knowledge integration, spatial understanding, and other aspects. We prompt GPT4-V to generate Q&As in six different aspects based on images captured from four different views of a 3D subject, as illustrated in Fig. 3.
Figure 4: 3D MM-Vet dataset sample. A wealth of precise evaluation metrics enable a comprehensive assessment.
Embodied Object Understanding  A comprehensive understanding of the spatial positions and semantics at the part level is crucial to facilitate effective object grasping and interaction in embodied scenarios. Fortunately, the GAPartNet [47] provides rich part annotations, including semantics and poses, which are instrumental in constructing instruction-tuning data for embodied interactive parts of a subject. Specifically, given a 3D object, questions are formulated based on the semantics of its different parts, and answers are constructed in both the semantics and 3D positions. The positions are represented as 6-DoF 3D bounding boxes in a straightened Python multidimensional list format, denoted as [[x1, y1, z1], [x2, y2, z2], …, [x8, y8, z8]], to meet characteristics of the textual dialogues response in LLMs. The canonical space of the object determines the sequence of coordinates. Using bounding box coordinates leverages the inherent spatial relationship, allowing LLMs to readily learn these patterns and generate accurate output coordinates. This approach can offer specific position information for embodied manipulation.
@@ 2.3 ReCon++: Scaling Up 3D Representation Learning
However, existing 3D cross-modal representation learning methods [170, 94] mainly distill high-resolution object features from single-view 2D foundation models, resulting in a unilateral shape understanding. Besides, they generally employ multi-view images as a data augmentation strategy, imposing the learned representation to the average representation of all views. Thus, the accurate 3D shape information is missing. Recently, ReCon [129] utilizes contrast guided by reconstruction to address the pattern disparities between local masked data modeling and global cross-modal alignment. This results in remarkable performance in various tasks, including transfer learning, zero-shot classification, and part segmentation. However, its potential is hindered by the scarcity of pretraining data [13].
@@
Denote $N$ as the number of multi-view images, $I_{i}$ is the image feature from $i$-th view, and $Q_{i}$ represents the global query of $i$-th view. Following Carion et al., we search for an optimal permutation $\sigma$ of $N$ elements with the lowest cost:
@@
3 3D MM-Vet: 3D Multimodal Comprehension Evaluation Benchmark
@@
    Table 1: Fine-tuned 3D recognition on ScanObjectNN and ModelNet40. BG, ON, and RS are short for OBJ_BG, OBJ_ONLY, and PB_T50_RS, respectively. Overall accuracy (%) with voting [98] is reported. ${}^{\dagger}$: results with a post-pretraining stage [18].
    Method ScanObjectNN ModelNet40
    BG ON RS 1k P 8k P
    Supervised Learning Only
    PointNet [126] 73.3 79.2 68.0 89.2 90.8
    PointNet++ [127] 82.3 84.3 77.9 90.7 91.9
    DGCNN [157] 82.8 86.2 78.1 92.9 -
    PointMLP [108] - - 85.4 94.5 -
    PointNeXt [131] - - 87.7 94.0 -
    Transformer [153] 83.04 84.06 79.11 91.4 91.8
    with Self-Supervised Representation Learning
    Point-BERT [180] 87.43 88.12 83.07 93.2 93.8
    Point-MAE [119] 90.02 88.29 85.18 93.8 94.0
    Point-M2AE [185] 91.22 88.81 86.43 94.0 -
    ACT [36] 93.29 91.91 88.21 93.7 94.0
    TAP  [158] - - 88.5 94.0 -
    VPP [130] 93.11 91.91 89.28 94.1 94.3
    I2P-MAE [188] 94.15 91.57 90.11 94.1 -
    ULIP-2 [170] - - 91.5 - -
    ReCon [129] 95.35 93.80 91.26 94.5 94.7
    ${}^{\dagger}$ [18] 95.8 95.2 91.9 94.4 94.6
    ${}^{\dagger}$ [18] 97.2 96.6 93.4 94.7 94.9
    ${}^{\dagger}$ 98.62 96.21 93.34 94.6 94.8
    ${}^{\dagger}$ 98.80 97.59 95.25 94.8 95.0
@@
    Embodied Interaction: The utilization scope of multimodal LLMs extends into the field of embodied interaction, facilitated by the utilization of instruction-following data. Our evaluation system tests their capacity by formally requesting LLMs to provide execution steps toward an instruction. This approach aims to establish connections for handling Embodied Interaction tasks [70, 39].
To prevent any overlap with training data, our collection of 3D models is sourced exclusively from Turbosquid [142], a platform not included in the acquisition lists of Objaverse [29] and ShapeNet [13]. We meticulously curated a dataset of 59 3D models, generating 232 Q&As for evaluation purposes. In our pursuit of a precise assessment of single-task capabilities, each question is designed to test only one specific capacity outlined earlier. Every question is paired with a corresponding answer tailored to the particular 3D model, serving as the ground truth. More details and analysis can be found in Appendix B.
Table 2: Zero-shot 3D recognition on Objaverse-LVIS [29], ModelNet40 [164] and ScanObjectNN [151]. Ensembled: pretraining with four datasets, Objaverse [29], ShapeNet [13], ABO [22] and 3D-FUTURE [42], following OpenShape [94]. ${}^{\dagger}$: Uni3D employs an EVA-CLIP-E [146] teacher that is larger and improves performance, while other methods employ OpenCLIP-bigG [74].
@@ 4 Experiments
Table 3: Zero-shot 3D multimodal comprehension evaluation of core VL capabilities in 3D context on 3D MM-Vet. Rec: General Visual Recognition, Know: Knowledge, Gen: Language Generation, Spat: Spatial Awareness, Emb: Embodied Interaction.
@@
LLaVA-13B [93] 1-View Img. 40.0 55.3 51.3 43.2 51.1 47.9
DreamLLM-7B [37] 4-View Img. 42.2 54.4 50.8 48.9 54.5 50.3
GPT-4V [115] 1-View Img. 53.7 59.5 61.1 54.7 59.0 57.4
GPT-4V [115] 4-View Img. 65.1 69.1 61.4 52.9 65.5 63.4
PointBind&LLM [54] Point Cloud 16.9 13.0 18.5 32.9 40.4 23.5
PointLLM-7B [166] Point Cloud 40.6 49.5 34.3 29.1 48.7 41.2
PointLLM-13B [166] Point Cloud 46.6 48.3 38.8 45.2 50.9 46.6
ShapeLLM-7B Point Cloud 45.7 42.7 43.4 39.9 64.5 47.4
ShapeLLM-13B Point Cloud 46.8 53.0 53.9 45.3 68.4 53.1
@@ 4.2 Multimodal Comprehension with ShapeLLM
Quantitative Analysis  To assess the comprehensive capabilities of ShapeLLM, we first quantitatively compare various baselines and our model on the proposed 3D MM-Vet benchmark using GPT-4. Following ModelNet-C [136] and ModelNet40-C [144], we construct 3D MM-Vet-C to benchmark the robustness against 3D corruptions.
  • 3D MM-Vet. Tab. 3 shows the detailed results of ShapeLLM on different tasks of 3D MM-Vet. It can be observed that ShapeLLM significantly outperforms PointLLM [166] across various metrics, particularly in Embodied Tasks. This substantiates our model’s versatile capability in addressing real-world scenario tasks.
  • 3D MM-Vet-C. Tab. 4 shows the comparison of model robustness against “single-view”, “jitter” and “rotate” corruptions, which are the most common corruptions in real-world scenarios. The results demonstrate significantly superior robustness of ShapeLLM against corruption, indicating stronger potential in real-world applicability.
Qualitative Analysis  Fig. 5 illustrates qualitative examples of ShapeLLM in multimodal dialogue. ShapeLLM is capable of supporting general VQA, embodied task and action planning, as well as 6-DoF pose estimation. Notably, due to the strict spatial relationship inherent in 6-DoF bounding box coordinates, we observe that LLMs easily grasp such patterns and consistently produce valid coordinates.
Table 4: Zero-shot 3D multimodal comprehension evaluation of robustness on 3D MM-Vet-C. Clean: no corruptions. Single-View: randomly select a camera viewpoint within the unit sphere and generate a single viewpoint within the FoV on polar coordinates. Jitter: Gaussian jittering with noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ and $\sigma=0.01$. Rotate: random SO(3) rotation sampling over X-Y-Z Euler angle $(\alpha,\beta,\gamma)\sim\mathcal{U}(-\theta,\theta)$ and $\theta=\pi/6$.

### v3

# https://qizekun.github.io/shapellm/

@@ Abstract
This paper presents ShapeLLM, the first 3D Multimodal Large Language Model (LLM) designed for embodied interaction, exploring a universal 3D object understanding with 3D point clouds and languages. ShapeLLM is built upon an improved 3D encoder by extending ReCon [135] to ReCon++ that benefits from multi-view image distillation for enhanced geometry understanding. By utilizing ReCon++ as the 3D point cloud input encoder for LLMs, ShapeLLM is trained on constructed instruction-following data and tested on our newly human-curated benchmark, 3D MM-Vet. ReCon++ and ShapeLLM achieve state-of-the-art performance in 3D geometry understanding and language-unified 3D interaction tasks, such as embodied visual grounding.
Keywords:
3D Point Clouds Large Language Models Embodied Intelligence 3D Representation Learning Zero-shot Learning
Figure 1: Demonstrations of ShapeLLM and ReCon++. We present ShapeLLM, the first 3D LLM designed for embodied interaction and spatial intelligence.
@@ 1 Introduction
    Selective Multi-View Distillation  Interacting with objects typically necessitates an intricate 3D understanding that involves knowledge at various levels and granularities. For instance, a whole-part high-level semantic understanding is needed for interactions like opening a large cabinet, while detailed, high-resolution (i.e., low-level) semantics are crucial for smaller objects like manipulating a drawer handle [181]. However, existing works mainly distill single-view high-resolution object features from 2D foundation models [138], providing a complementary understanding [37, 135, 175]. The potential of multi-view images, which offer abundant multi-level features due to view variation and geometry consistency [9, 149, 82, 61, 103, 66], is often neglected. ShapeLLM extends ReCon [135] to ReCon++  as the 3D encoder by integrating multi-view distillation. To enable the model to selectively distill views that enhance optimization and generalization, inspired by DETR [12], ReCon++  is optimized through adaptive selective matching using the Hungarian algorithm [85].
@@
    3D Visual Instruction Tuning  Instruction tuning has been proven effective in improving LLMs’ alignment capability [122, 126]. To realize various 3D understanding tasks with a universal language interface, ShapeLLM is trained through instruction-following tuning on constructed language-output data. However, similar to 2D visual instruction tuning [96, 4], the data-desert issue [37] is even worse since no object-level VQA data is available, unlike 2D [95]. To validate the efficacy of ShapeLLM, we first construct $\sim$45K instruction-following data using the advanced GPT-4V(ision) [120] on the processed Objaverse dataset [30] and 30K embodied part understanding data from GAPartNet [50] for supervised fine-tuning. Following MM-Vet [185], we further develop a novel evaluation benchmark named 3D MM-Vet. This benchmark is designed to assess the core vision-language capabilities, including embodied interaction in a 3D context, thereby stimulating future research. The 3D MM-Vet benchmark comprises 59 diverse Internet [^1] URL & License. 3D objects and 232 human-written question-answer pairs.
@@ 2 ShapeLLM
In this section, we first introduce the overall architecture of ShapeLLM. Then, we delve into two critical challenges faced in interactive 3D understanding: data desert [37] and representation of 3D point clouds. We present the detailed design of our method to tackle these challenges, respectively.
@@ 2.1 Overall Architecture
Figure 3: Qualitative visualization of the instruction-following and 3D MM-Vet data.
2.2 How to alleviate interactive 3D understanding Data Desert?
Most published 3D data is typically presented as 3D object-caption pairs, lacking an interactive style. Although a few concurrent works [172, 65] have attempted to construct interactive 3D understanding datasets, the questions-and-answers (Q&As) are primarily based on annotated captions, often providing a limited perspective without sufficient details. Additionally, those works have generally been limited to semantic understanding without considering embodied interaction. To address these limitations, our work constructs question-and-answer pairs based on multi-view images of a 3D object using GPT-4V(ision) [120]. For data diversity, we explicitly introduce six aspects as prompts, as illustrated Fig. 3(a). In the following, we provide the details about data collection and construction regarding general semantic understanding and embodied object understanding, respectively.
@@ 2.2 How to alleviate interactive 3D understanding Data Desert?
General Semantic Understanding  This aims to enhance the model’s generalization abilities in visual recognition, knowledge integration, spatial understanding, and other aspects. We prompt GPT4-V to generate Q&As in six different aspects based on images captured from four different views, as illustrated in Fig. 3(a).
Figure 4: Qualitative examples of the embodied interaction data.
Embodied Object Understanding  A comprehensive understanding of the spatial positions and semantics at the part level is crucial to facilitate effective object grasping and interaction in embodied scenarios. Fortunately, the GAPartNet [50] provides rich part annotations, including semantics and poses, which are instrumental in constructing instruction-tuning data for embodied interactive parts of a subject. Specifically, given a 3D object, questions are formulated based on the semantics of its different parts, and answers are constructed in both the semantics and 3D positions. The positions are represented as 6-DoF 3D bounding boxes in a straightened Python multidimensional list format, denoted as [[x1, y1, z1], [x2, y2, z2], …, [x8, y8, z8]], to meet characteristics of the textual dialogues response in LLMs. The canonical space of the object determines the sequence of coordinates. Using bounding box coordinates leverages the inherent spatial relationship, allowing LLMs to readily learn these patterns and generate accurate output coordinates. This approach can offer specific position information for embodied manipulation, as illustrated in Fig. 4.
@@ 2.3 ReCon++: Scaling Up 3D Representation Learning
However, existing 3D cross-modal representation learning methods [176, 97] mainly distill high-resolution object features from single-view 2D foundation models, resulting in a unilateral shape understanding. Besides, they generally employ multi-view images as data augmentation, imposing the learned representation to the average representation of all views. Thus, the accurate 3D shape information is missing. Recently, ReCon [135] utilizes contrast guided by reconstruction to address the pattern disparities between local masked data modeling and global cross-modal alignment. This results in remarkable performance in various tasks, including transfer learning, zero-shot classification, and part segmentation. However, its potential is hindered by the scarcity of pretraining data [13].
@@
Denote $N$ as the number of multi-view images, $I_{i}$ is the image feature from $i$-th view, and $Q_{i}$ represents the global query of $i$-th view. Following DETR [12], we search for an optimal permutation $\sigma$ of $N$ elements with the lowest cost:
@@
3 3D MM-Vet: Benchmarking 3D Comprehension
@@
Table 1: Fine-tuned 3D recognition on ScanObjectNN and ModelNet40. Overall accuracy (%) with voting [101] is reported. : Results with a post-pretraining stage [18].
Method ScanObjectNN ModelNet40
OBJ_BG OBJ_ONLY PB_T50_RS 1k P 8k P
Supervised Learning Only
PointNet [132] 73.3 79.2 68.0 89.2 90.8
PointNet++ [133] 82.3 84.3 77.9 90.7 91.9
DGCNN [163] 82.8 86.2 78.1 92.9 -
PointMLP [112] - - 85.4 94.5 -
PointNeXt [137] - - 87.7 94.0 -
with Self-Supervised Representation Learning
Point-BERT [186] 87.43 88.12 83.07 93.2 93.8
Point-MAE [124] 90.02 88.29 85.18 93.8 94.0
Point-M2AE [192] 91.22 88.81 86.43 94.0 -
Point2Vec [187] 91.2 90.4 87.5 94.8 -
ACT [37] 93.29 91.91 88.21 93.7 94.0
TAP  [164] - - 88.5 94.0 -
VPP [136] 93.11 91.91 89.28 94.1 94.3
I2P-MAE [195] 94.15 91.57 90.11 94.1 -
ULIP-2 [176] - - 91.5 - -
ReCon [135] 95.35 93.80 91.26 94.5 94.7
PointGPT-B [18] 95.8 95.2 91.9 94.4 94.6
PointGPT-L [18] 97.2 96.6 93.4 94.7 94.9
ReCon++-B 98.62 96.21 93.34 94.6 94.8
ReCon++-L 98.80 97.59 95.25 94.8 95.0
@@
    Embodied Interaction: The utilization scope of MLLMs extends into the field of embodied interaction, facilitated by the utilization of instruction-following data. Our evaluation system tests their capacity by formally requesting LLMs to provide execution steps toward an instruction. This approach aims to establish connections for handling Embodied Interaction tasks [73, 40].
Table 2: Zero-shot 3D recognition on Objaverse-LVIS [30], ModelNet40 [170] and ScanObjectNN [157]. Ensembled [97]: pretraining with four datasets, Objaverse [30], ShapeNet [13], ABO [23] and 3D-FUTURE [44]. : Uni3D employs a larger EVA-CLIP-E [152] teacher, while other methods employ OpenCLIP-bigG [77].
@@
To prevent any overlap with training data, our collection of 3D models is sourced exclusively from Turbosquid [148], a platform not included in the acquisition lists of Objaverse [30] and ShapeNet [13]. We meticulously curated a dataset of 59 3D models, generating 232 Q&As for evaluation purposes. In our pursuit of a precise assessment of single-task capabilities, each question is designed to test only one specific capacity outlined earlier. Every question is paired with a corresponding answer tailored to the particular 3D model, serving as the ground truth. Dataset samples are illustrated in Fig. 3(b). More details and analysis can be found in the supplemental material.
@@ 4 Experiments
Table 3: Zero-shot 3D multimodal comprehension of core VL capabilities in 3D context on 3D MM-Vet. Rec: General Visual Recognition, Know: Knowledge, Gen: Language Generation, Spat: Spatial Awareness, Emb: Embodied Interaction.
@@
LLaVA-13B [96] 1-View 2D Image 40.0 55.3 51.3 43.2 51.1 47.9
DreamLLM-7B [36] 4-View 2D Image 42.2 54.4 50.8 48.9 54.5 50.3
GPT-4V [120] 1-View 2D Image 53.7 59.5 61.1 54.7 59.0 57.4
GPT-4V [120] 4-View 2D Image 65.1 69.1 61.4 52.9 65.5 63.4
PointBind&LLM [57] 3D Point Cloud 16.9 13.0 18.5 32.9 40.4 23.5
PointLLM-7B [172] 3D Point Cloud 40.6 49.5 34.3 29.1 48.7 41.2
PointLLM-13B [172] 3D Point Cloud 46.6 48.3 38.8 45.2 50.9 46.6
ShapeLLM-7B 3D Point Cloud 45.7 42.7 43.4 39.9 64.5 47.4
ShapeLLM-13B 3D Point Cloud 46.8 53.0 53.9 45.3 68.4 53.1
@@ 4.2 Multimodal Comprehension with ShapeLLM
Quantitative Analysis  To assess the comprehensive capabilities of ShapeLLM, we first quantitatively compare various baselines and our model on the proposed 3D MM-Vet using GPT-4. Following ModelNet-C [142] and ModelNet40-C [150], we construct 3D MM-Vet-C to benchmark the robustness against 3D corruptions.
Table 4: Zero-shot 3D multimodal comprehension of robustness on 3D MM-Vet-C. Clean: no corruptions. Single-View: randomly select a camera viewpoint within the unit sphere and generate a single viewpoint within the FoV on polar coordinates. Jitter: Gaussian jittering with noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ and $\sigma=0.01$. Rotate: random SO(3) rotation sampling over X-Y-Z Euler angle $(\alpha,\beta,\gamma)\sim\mathcal{U}(-\theta,\theta)$ and $\theta=\pi/6$.
@@
i) 3D MM-Vet.  Tab. 3 shows the detailed results of ShapeLLM on different tasks of 3D MM-Vet. It is observed that ShapeLLM significantly outperforms PointLLM [172] across various metrics, particularly in Embodied Tasks. This substantiates our model’s versatile capability in addressing real-world tasks.
ii) 3D MM-Vet-C.  Following the ModelNet-C [142] and ModelNet40-C [150], we construct 3D MM-Vet-C to benchmark the robustness against 3D corruptions. Tab. 4 shows the comparison of robustness against “single-view”, “jitter”, and “rotate” corruptions, which are the most common in real scenarios. The “single-view” issue is the most critical challenge since obtaining the complete point clouds is non-trivial, similar to multi-view images. Therefore, everyday real-world robots only get single-view 3D perceptions with sensors such as RGB-D [59]. The results demonstrate significantly superior robustness of ShapeLLM, indicating stronger potential in real-world applicability.
Baseline Improvement  Can we improve the baseline to bridge the gap between PointLLM and ShapeLLM? In Tab. 5, we study two technical factors that are contributed by ShapeLLM: 3D point cloud encoder and SFT data.
Figure 6: 3D multimodal dialogue using single-view point cloud inputs. All answers are generated by ShapeLLM-13B with single-view occluded inputs. ShapeLLM achieves outstanding robustness against such commonly met occlusion in the real world.
Table 5: Ablation study on baseline improvements. Results are tested on 3D MM-Vet with the baseline model PointLLM-13B [172] using different point encoders and SFT data.
Encoder SFT Data Rec Know Gen Spat Emb Total
ULIP-2 [176] PointLLM 46.6 48.3 38.8 45.2 50.9 46.6
ReCon++ PointLLM 47.5 52.8 43.6 44.9 54.5 50.8
ReCon++ Ours 46.8 53.0 53.9 45.3 68.4 53.1
i) Improvement from encoder.
First, by changing PointLLM’s encoder to ReCon++, a significant improvement of +4.20% is obtained. This demonstrates the significantly better 3D representation extraction of ReCon++ compared to ULIP-2. It is consistent with previous findings in Tab. 1 and Tab. 2 that ReCon++ outperforms ULIP-2 by a large margin regarding 3D representation transferring learning and zero-shot learning.
ii) Improvement from data.  As stated in Sec. 2.2, we have constructed instruction-following data for supervised fine-tuning (SFT) using GPT-4V involving diverse topics. By further using the SFT data curated by us, PointLLM’s performance gap to ShapeLLM has been fulfilled. This demonstrates the superiority of our SFT data, where the decent quality comes from the advanced GPT4-V using multi-view images and the topics covered in the data.
Qualitative Analysis  Fig. 5 illustrates qualitative examples of ShapeLLM in multimodal dialogue. ShapeLLM can support general VQA, embodied task and action planning, and 6-DoF pose estimation. Notably, LLMs easily grasp such patterns and consistently produce valid coordinates due to the strict spatial relationship inherent in 6-DoF bounding box coordinates. Fig. 6 shows the examples of ShapeLLM-13B’s response using single-view point cloud inputs, demonstrating surprisingly outstanding robustness in processing such occlusion. This is crucial for the practical deployment of real machines, as single-view point clouds can be easily obtained from RGB-D cameras.